OpenStudy (kailanineon):

Solve and express the solution set in simplest form. 8x-2/9 = 2/1 9/4 5/2 9/2 9/8

OpenStudy (anonymous):

does it really say \(\frac{2}{1}\) ?

OpenStudy (kailanineon):

Yes

OpenStudy (anonymous):

did they not realize that \(\frac{2}{1}=2\) ?

OpenStudy (the_fizicx99):

Lol xD ^

OpenStudy (kailanineon):

Apparently not.

OpenStudy (anonymous):

next question is it \[\frac{8x-9}{2}=2\] or is it \[8x-\frac{9}{2}=2\]?

OpenStudy (kailanineon):

The first one, just 8x-2 over 9

OpenStudy (anonymous):

multiply both sides by \(2\) and get \[8x-9=18\] then solve that one it two steps

OpenStudy (anonymous):

steps are 1) add 9 2) divide by 8